### v2

@@ Abstract
Aims. In this paper, we present and discuss the adaptation of state-of-the-art algorithms, initially designed for ADI, to jointly leverage angular and reference-star differential imaging (ARDI) for direct high-contrast imaging of circumstellar disks.
Methods. Using a collection of high-contrast imaging data sets, we assess the performance of ARDI in comparison to ADI and RDI based on iterative principal component analysis (IPCA). These diverse data sets are acquired under various observing conditions and include the injection of synthetic disk models at various contrast levels. We also considered reference stars with different levels of correlation with the science targets.
Results. Our results demonstrate that ARDI with IPCA improves the quality of recovered disk images and the sensitivity to planets embedded in disks, compared to ADI or RDI individually. This enhancement is particularly pronounced when dealing with extended sources exhibiting highly ambiguous structures that cannot be accurately retrieved using ADI alone, and when the quality of the reference frames is suboptimal, leading to an underperformance of RDI. We finally apply our method to a sample of real observations of protoplanetary disks taken in star-hopping mode, and propose to revisit the protoplanetary claims associated with these disks. Among eight proposed protoplanets claimed through velocity kinks or direct imaging, none of them were re-detected in our new processed images.
@@ 1 Introduction
High-contrast imaging (HCI) is a technique that aims to observe faint signals near bright sources, such as exoplanets and circumstellar disks. After the majority of the starlight and aberrations caused by atmospheric turbulence have been removed by the coronagraph and the adaptive optics respectively, residual atmospheric speckles and quasi-static speckles not seen by the AO system remain (Racine et al. 1999). These speckles must be removed to reveal the faint circumstellar signal hiding in the vicinity of the star. Several observing strategies aim to provide a lever to distinguish the speckle field from the disk and planet signal, such as spectral differential imaging (SDI, Racine et al. 1999; Marois et al. 2000) and polarimetric differential imaging (PDI, Kuhn et al. 2001). In this manuscript, we will focus on two strategies.Firstly, angular differential imaging (ADI, Marois et al. 2006) consists of capturing images over time while targeting a star and fixing the orientation of the telescope pupil, causing the field of view to rotate due to the Earth’s rotation. The discriminant employed by the ADI strategy to differentiate the speckle from actual signals is that, unlike the field of view that rotates according to the parallactic angle, the speckle field resulting from imperfections in the instrument will remain mostly fixed. Angular diversity is typically leveraged by discarding the static component (which represent the speckle field) to retrieve the rotating circumstellar signal. Secondly, reference-star differential imaging (RDI, Mawet et al. 2009; Lafrenière et al. 2009; Ruane et al. 2019) involves using observations of reference star(s) which share similar characteristics with the scientific target of interest, either taken during data acquisition or selected from an archive library. The primary goal is to acquire reference images that exhibit a similar speckle pattern. References are leveraged as a model of the speckle field, to be subtracted from the observed data after being appropriately scaled in flux.
@@
The choice of an appropriate observing strategy depends on the specific scientific goals, the nature of the target being observed, and the quality of available resources. Each of these high-contrast imaging strategies has its strengths and limitations. ADI is effective for distinguishing faint point-like sources, but does require a sufficient parallactic angle rotation to function correctly and can cause deformations in extended sources due to rotational invariance (Milli et al. 2012). These deformations depend on the amount of rotation and the corresponding degree of rotational symmetry, considering the disk morphology (Juillard et al. 2023). This limitation not only hampers the ability to capture the most face-on disks, but more generally applies to any extended source morphology, leading to challenging-to-correct deformations that can mislead the interpretation of the astronomer. Meanwhile, RDI is efficient for stable speckle patterns and can produce high-fidelity images of disks, but generally requires dedicated observations of an appropriate reference star in very similar observing conditions. It is less effective in the case of particularly challenging data sets, such as faint disks, time-varying speckle patterns, or when no well-suited reference star is available, which can be the case for archival ADI data sets.
@@
To best utilise the capabilities of each strategy, employing a suitable post-processing algorithm is required. Principal component analysis (PCA, Soummer et al. 2012; Amara & Quanz 2012) is widely used and can be applied to both ADI and RDI, but can suffer from overly aggressive PSF subtraction, resulting in deformations of the circumstellar signal (Milli et al. 2012). In the case of ADI, PCA and other PSF-subtraction algorithms such as median-subtraction (c-ADI, Marois et al. 2006) and LOCI (locally optimised combination of images, Lafrenière et al. 2007) work well for point-like sources but generally struggle with extended structures like disks. Despite many efforts to compensate for these limitations, more recent algorithms such as iterative PCA (IPCA Pairet et al. 2021; Stapper & Ginski 2022) or inverse problem (IP) approaches (MAYONNAISE, REXPACO, MUSTARD, Pairet et al. 2021; Flasseur et al. 2021; Juillard et al. 2022) are still prone to deformations of extended structures, in particular when these structures include signals that appear static throughout the ADI image sequence. This means that angular diversity is not a suitable discriminator to distinguish such signal from the quasi-static speckle field. Regarding RDI, other algorithms such as data imputation with sequential non-negative matrix factorization (DI-sNMF, Ren et al. 2018) and the Karhunen-Loève transform with data imputation (DIKL, Ren 2023) have demonstrated efficiency in optimally extracting circumstellar signals and mitigating the risk of over-subtraction. While RDI can be highly effective, it necessitates that the speckle pattern remains relatively stable over time, and requires sufficiently correlated reference stars to produce high fidelity images.
Table 1: Notations
@@
Combining diverse observing strategies might assist in mitigating the distinct weaknesses of each approach, as exemplified by Lawson et al. (2022), who combined RDI with PDI, or by Wahhaj et al. (2013), Christiaens et al. (2019) and Flasseur et al. (2022), who merged ADI with SDI (ASDI) for extended sources. Previous studies in the scientific literature have explored the simultaneous use of RDI and ADI. These attempts typically involve using techniques such as median-subtraction or PCA, and applying them to the concatenation of the reference and science cubes (e.g., Carter et al. 2023, Wallack et al. in press). However, there has not been a dedicated effort to identify the most effective method of combining ADI and RDI, especially in the context of disk imaging. Previous research in this context was rather focused on comparing the performance of ADI with RDI (Ruane et al. 2019; Xie et al. 2022). In response to the described challenges and inspired by the many efforts made in the past few years to develop efficient novel algorithms to retrieve extended sources using ADI and RDI strategies separately, this paper aims to explore how to adapt the IPCA algorithm to jointly leverage angular and reference-star differential imaging (ARDI) in a simultaneous approach. In Sect. 2, we explain the theoretical foundation and limitations of state-of-the-art IPCA algorithms employing ADI and RDI separately. In Sect. 3, we adapt the IPCA algorithm to jointly leverage ADI and RDI. We also explore the question of parameter optimization. In Sect. 4, we conduct a comprehensive comparative evaluation of ARDI against RDI and ADI, all using an IPCA-based algorithm, and assess their performance across diverse scenarios. In Sect. 5, we test IPCA with ARDI on a sample of real observations of protoplanetary disks, and reevaluate the planet candidates claimed in these disks.
Notations. In the rest of the manuscript, we use the following notations, which are shared with a previous publication (Juillard et al. 2023): $Y\in\mathbb{R}^{n,m\times m}$ denotes the ADI sequence (also known as the ADI cube), with $n$ being the number of frames, and where each frame is a vectorized square image of size $m\times m$ pixels. The ADI sequence is composed of the circumstellar signal $d$ and the speckle field $S$. The estimate of a parameter is written with an over bar (e.g., $\bar{S}$). We use $\bar{S}\in\mathbb{R}^{n,m\times m}$ to denote the cube of the estimated speckle field, which is composed of $n$ unique frames. We use $\bar{d}\in\mathbb{R}^{+,m\times m}$ to denote the estimation of the circumstellar signal, which is one single positive image, commonly shared by all frames of the cube. Finally, $L_{r}\in\mathrm{R}^{r,m\times m}$ denotes the library of reference stars containing $r$ frames. Matrix products are written using a dot symbol ($\cdot$). A complete list of all the notation used in this manuscript is presented in Table 1.
@@ 2 IPCA algorithms for ADI and RDI processing
Iterative PCA has demonstrated proficiency in preserving disk signal when leveraging angular diversity (Pairet et al. 2021; Stapper & Ginski 2022; Juillard et al. 2023). The iterative aspect enables the usage of a model that constrains the positivity of the estimates to enhance the performance of PCA. In this section, we re-introduce the IPCA method using either ADI or RDI. These approaches will enable us to establish, in Sect. 4, a consistent comparison with our proposed adaption of IPCA for ARDI, as all three strategies, or combinations of strategies, make use of an IPCA-based algorithm. Alongside the presentation of these algorithms, we propose an explanation of the concept of a fixed point, which is fundamental to IPCA algorithms. Understanding the concept of a fixed point is not only essential for these algorithms, but also crucial for appreciating parameter optimization.
@@ 2.1 ADI with IPCA
Figure 1: Best disk image, according to the SSIM metric, obtained with the IPCA algorithm leveraging ADI-only while processing a synthetic data set (disk #B, cube #3 at a contrast of $10^{-3}$; see Section 4). The image on the left displays different apertures placed on (1) the brightest part of the disk, (2) a region of the disk that is rotation invariant, (3) a noise region, and (4) the region near the coronagraphic region where the variation of the speckle field is higher. On the right-hand side, we plot the evolution of the integrated flux computed in these different apertures through the IPCA iterations. The red vertical line indicates the position of the best image among all iterations according to SSIM. Each curve in the figure is normalized by its mean. The x-axis represents each iteration of the IPCA process, starting at rank 1, up to rank 10, with 10 iterations per rank. The gray horizontal line, placed at $y=1$, indicates the position where the flux is equal to its mean.
@@
Figure 2: Same as Fig. 1, but for RDI-only IPCA. The set of reference frames used here is presented in Sect. 4 and is referred to as “optimal”. The x-axis represents each iteration of the IPCA process, for 10 iterations using a rank-1 PCA.
In Fig. 1, we show the evolution of the disk estimate through IPCA iterations on an example synthetic data set. The IPCA is parameterized such that it performs 10 iterations per rank, while starting at rank $q=1$ and increasing the rank up to $q=10$. The image on the left provides the best estimation of the injected disk according to the structural similarity index measure (SSIM, Wang et al. 2004). This metric, which we also use in Section 4, is particularly adapted for assessing image similarity. It evaluates not only the pixel-wise differences between the images but also considers structural information such as contrast, luminance, and structure on various windows of the two images. Indeed, ranking disk image estimations is a multifactorial problem that often necessitates finding a satisfactory compromise between fewer geometrical biases and less noise residuals or errors in the overall restored flux.
@@ 2.2 RDI with IPCA
The implementations of the RDI algorithm found in the scientific literature (Schneider et al. 2009; Soummer et al. 2012; Ren et al. 2018) involves the subtraction of an estimated speckle pattern based on a scaled version of the reference images. We propose an adapted version of the IPCA algorithm that employs an iterative process to optimize the projection of the estimated reference. For a given set of references $L_{r}\in\mathbb{R}^{r,m\times m}$ of $r$ reference images, we will refer to ${V}^{\rm ref}_{q}\in\mathbb{R}^{q,m\times m}$ as the collection of $q$ PCs estimated by PCA on the reference library. We define the operator $\mathcal{H}_{q}^{L_{r}}(Y)$, which, similarly to $\mathcal{H}_{q}(Y)$, returns $\bar{S}_{i+1}$ by projecting the science cube into the principal components, which are now computed on the reference library. Note that $\mathcal{H}_{q}(Y)=\mathcal{H}_{q}^{Y}(Y)$. The subsequent iterative process can then be expressed as:
@@
We display in Fig. 2 the convergence process to find the optimal projection while using IPCA-RDI of rank $q=1$. This example was produced using the same test dataset as in Fig. 1. In this example, we can observe that the flux exhibits a very clear convergence, which is achieved in only a few iterations. This iterative process enables to mitigate the over-subtraction effect, and significantly improves the results compared to classical PCA. RDI however remains intrinsically sensitive to the stability of the speckle pattern over time, and necessitates the availability of sufficiently correlated reference stars.
@@ 3 Combining angular and reference-star differential imaging (ARDI)
Simultaneously employing both RDI and ADI strategies is expected to help mitigate the individual limitations of each strategy. The combination of RDI with ADI holds the potential of offering several advantages in the field of high-contrast imaging for exoplanet and disk detection around stars. Here, our primary focus is to adapt the IPCA approach, initially designed for ADI, to leverage ARDI. Other advanced algorithms using both ADI and RDI, including IP-based approaches, could also be considered, but are beyond the scope of this paper. In this section, we will also explore the problem of parameter optimization for IPCA.
@@
where the operation [_;_]__[\_\;;\;\_][ _ ; _ ] represents the concatenation of two collections of images, creating a set of images ($L_{n+r,i}$) with a size of $\mathbb{R}^{n+r,m\times m}$ that contains both the data and reference images. The ratio between the size $r$ of the reference library and the number $n$ of frames in the ADI cube must be chosen carefully for optimal results. We will refer to $r/n$ as the reference frame ratio. The proposed algorithm can be classified as semi-supervised, as it simultaneously leverages reference frames for the speckle field (which can be regarded as labeled data), and the actual (unlabeled) data to provide an optimal estimation.
Figure 4: Evolution of the IPCA with ARDI estimate through the iterations, using the same data set as in Fig. 3. The blue curve relates to the left y-axis and displays the value of the SSIM for each estimate. The orange curve relates to the right y-axis and represents the gradient computed over the full frame, as described in Eq. 4. The $g_{\rm tol}$ has been set to $10^{-2}$ and is represented on the figure by the two grey lines. The black dot indicates the iteration where the gradient curve reaches the $g_{\rm tol}$ and does not exceed it in further steps. The x-axis represents each iteration, where the minor ticks represent a single iteration, and the major ticks represent a rank update.
In Fig. 3, we monitor the evolution of the disk estimate through the IPCA iterations while using the ARDI strategy, similarly to the test presented in Fig. 1 with ADI only. We used identical parametrization (ten ranks, ten iterations per rank), and a reference frame ratio ($r/n$) of one. Conversely to what we observed with ADI only, three out of the four apertures seem to converge when using reference images. Results obtained with ARDI do not exhibit the deformation due to rotation invariant flux (located in aperture #2). However, the background noise in the final image increases throughout the iterations. Consequently, determining where to stop the iterative process involves choosing a good compromise between noise subtraction (non-propagation) and filtering/recovery of extended signals (convergence). A common criterion to assess convergence is setting a stopping criterion based on the gradient of the iterate. In our case, the criterion, based on the disk estimate, reads as follows:
@@
Figure 5: Pearson correlation coefficient of the selected reference frames computed individually and averaged over each frame of the raw ADI data before the injection of simulated disks or planets, for three different selections of references: (left) optimal references selected, (middle) a random sample of each cube’s dedicated references, including the optimal references, in even proportion, and (right) a random sample of references excluding the optimal references. These coefficients were calculated within an annulus with an outer radius of 60 pixels and inner radius of 30 pixels, and are classified from the most correlated to the least correlated. The text box below each figure displays the mean PCC ($\mu$) and the variance ($\sigma$) of the reference sample relative to its associated data cube.

### v3

@@ Abstract
Aims. In this paper, we present and discuss the adaptation of state-of-the-art algorithms, initially designed for ADI, to jointly leverage angular and reference-star differential imaging (ARDI) for the direct HCI of circumstellar disks.
Methods. Using a collection of HCI datasets, we assessed the performance of ARDI in comparison to ADI and RDI based on iterative principal component analysis (IPCA). These diverse datasets were acquired under various observing conditions and include the injection of synthetic disk models at various contrast levels. We also considered reference stars with different levels of correlation with the science targets.
Results. Our results demonstrate that ARDI with IPCA improves the quality of recovered disk images and the sensitivity to planets embedded in disks, compared to ADI or RDI individually. This enhancement is particularly pronounced when dealing with extended sources exhibiting highly ambiguous structures that cannot be accurately retrieved using ADI alone, and when the quality of the reference frames is suboptimal, leading to an underperformance of RDI. We finally applied our method to a sample of real observations of protoplanetary disks taken in star-hopping mode, and propose to revisit the protoplanetary claims associated with these disks. Among eight proposed protoplanets claimed through velocity kinks or direct imaging, none of them were re-detected in our new processed images.
@@ 1 Introduction
High-contrast imaging (HCI) is a technique that aims to observe faint signals near bright sources, such as exoplanets and circumstellar disks. After the majority of the starlight and aberrations caused by atmospheric turbulence have been removed by the coronagraph and the adaptive optics (AO), respectively, residual atmospheric speckles and quasi-static speckles not seen by the AO system remain (Racine et al. 1999). These speckles must be removed to reveal the faint circumstellar signal hiding in the vicinity of the star. Several observing strategies aim to provide a lever to distinguish the speckle field from the disk and planet signal, such as spectral differential imaging (SDI, Racine et al. 1999; Marois et al. 2000) and polarimetric differential imaging (PDI, Kuhn et al. 2001). In this paper, we focus on two strategies. Firstly, angular differential imaging (ADI, Marois et al. 2006) consists of capturing images over time while targeting a star and fixing the orientation of the telescope pupil, causing the field of view to rotate due to the Earth’s rotation. The discriminant employed by the ADI strategy to differentiate the speckle from actual signals is that, unlike the field of view that rotates according to the parallactic angle, the speckle field resulting from imperfections in the instrument remains mostly fixed. Angular diversity is typically leveraged by discarding the static component (which represents the speckle field) to retrieve the rotating circumstellar signal. Secondly, reference-star differential imaging (RDI, Mawet et al. 2009; Lafrenière et al. 2009; Ruane et al. 2019) involves using observations of reference star(s) that share similar characteristics with the scientific target of interest, either taken during data acquisition or selected from an archive library. The primary goal is to acquire reference images that exhibit a similar speckle pattern. References are leveraged as a model of the speckle field, which is to be subtracted from the observed data after being appropriately scaled in flux.
@@
The choice of an appropriate observing strategy depends on the specific scientific goals, the nature of the target being observed, and the quality of available resources. Each of these high-contrast imaging strategies has its strengths and limitations. ADI is effective at distinguishing faint point-like sources, but it does require a sufficient parallactic angle rotation to function correctly and can cause deformations in extended sources due to rotational invariance (Milli et al. 2012). These deformations depend on the amount of rotation and the corresponding degree of rotational symmetry, considering the disk morphology (Juillard et al. 2023). This limitation not only hampers most face-on disks from being captured, but more generally applies to any extended source morphology, leading to challenging-to-correct deformations that can mislead the interpretation of the astronomer. Meanwhile, RDI is efficient for stable speckle patterns and can produce high-fidelity images of disks, but generally requires dedicated observations of an appropriate reference star in very similar observing conditions. It is less effective in the case of particularly challenging datasets, such as faint disks, time-varying speckle patterns, or when no well-suited reference star is available, which can be the case for archival ADI datasets.
@@
To best utilize the capabilities of each strategy, employing a suitable post-processing algorithm is required. Principal component analysis (PCA, Soummer et al. 2012; Amara & Quanz 2012) is widely used and can be applied to both ADI and RDI, but it can suffer from overly aggressive PSF (point spread function) subtraction, resulting in deformations of the circumstellar signal (Milli et al. 2012). In the case of ADI, PCA and other PSF-subtraction algorithms such as median subtraction (c-ADI, Marois et al. 2006) and LOCI (locally optimized combination of images, Lafrenière et al. 2007) work well for point-like sources but generally struggle with extended structures such as disks. Despite many efforts to compensate for these limitations, more recent algorithms such as iterative PCA (IPCA Pairet et al. 2021; Stapper & Ginski 2022) or inverse problem (IP) approaches (MAYONNAISE, REXPACO, and MUSTARD; Pairet et al. 2021; Flasseur et al. 2021; Juillard et al. 2022) are still prone to deformations of extended structures, in particular when these structures include signals that appear static throughout the ADI image sequence. This means that angular diversity is not a suitable discriminator to distinguish this type of signal from the quasi-static speckle field. Regarding RDI, other algorithms such as data imputation with sequential non-negative matrix factorization (DI-sNMF, Ren et al. 2018) and the Karhunen-Loève transform with data imputation (DIKL, Ren 2023) have demonstrated efficiency in optimally extracting circumstellar signals and mitigating the risk of over-subtraction. While RDI can be highly effective, it necessitates that the speckle pattern remain relatively stable over time, and requires sufficiently correlated reference stars to produce high fidelity images.
Table 1: Notations.
@@
Combining diverse observing strategies might assist in mitigating the distinct weaknesses of each approach, as exemplified by Lawson et al. (2022), who combined RDI with PDI, or by Wahhaj et al. (2013), Christiaens et al. (2019), and Flasseur et al. (2022), who merged ADI with SDI (ASDI) for extended sources. Previous studies in the scientific literature have explored the simultaneous use of RDI and ADI. These attempts typically involve using techniques such as median subtraction or PCA, and applying them to the concatenation of the reference and science cubes (e.g., Carter et al. 2023, Wallack et al. in press). However, there has not been a dedicated effort to identify the most effective method of combining ADI and RDI, especially in the context of disk imaging. Previous research in this context was rather focused on comparing the performance of ADI with RDI (Ruane et al. 2019; Xie et al. 2022). In response to the described challenges and inspired by the many efforts made in the past few years to develop efficient novel algorithms to retrieve extended sources using ADI and RDI strategies separately, this paper aims to explore how one can adapt the IPCA algorithm to jointly leverage angular and reference-star differential imaging (ARDI) in a simultaneous approach. In Sect. 2, we explain the theoretical foundation and limitations of state-of-the-art IPCA algorithms employing ADI and RDI separately. In Sect. 3, we explain how we adapted the IPCA algorithm to jointly leverage ADI and RDI. We also explore the question of parameter optimization. For Sect. 4, we conducted a comprehensive comparative evaluation of ARDI against RDI and ADI, all using an IPCA-based algorithm, and assessed their performance across diverse scenarios. In Sect. 5, we explain how we tested IPCA with ARDI on a sample of real observations of protoplanetary disks, and reevaluated the planet candidates claimed in these disks.
Notations. In the rest of the paper, we use the following notations, which are shared with a previous publication (Juillard et al. 2023): $Y\in\mathbb{R}^{n,m\times m}$ denotes the ADI sequence (also known as the ADI cube), with $n$ being the number of frames, and where each frame is a vectorized square image of size $m\times m$ pixels. The ADI sequence is composed of the circumstellar signal $d$ and the speckle field $S$. The estimate of a parameter is written with an over bar (e.g., $\bar{S}$). We use $\bar{S}\in\mathbb{R}^{n,m\times m}$ to denote the cube of the estimated speckle field, which is composed of $n$ unique frames. We use $\bar{d}\in\mathbb{R}^{+,m\times m}$ to denote the estimation of the circumstellar signal, which is one single positive image, commonly shared by all frames of the cube. Finally, $L_{r}\in\mathrm{R}^{r,m\times m}$ denotes the library of reference stars containing $r$ frames. Matrix products are written using a dot symbol ($\cdot$). A complete list of all the notation used in this paper is presented in Table 1.
@@ 2 IPCA algorithms for ADI and RDI processing
Iterative PCA has demonstrated proficiency in preserving disk signal when leveraging angular diversity (Pairet et al. 2021; Stapper & Ginski 2022; Juillard et al. 2023). The iterative aspect enables the usage of a model that constrains the positivity of the estimates to enhance the performance of PCA. In this section we reintroduce the IPCA method using either ADI or RDI. These approaches will enable us to establish, in Sect. 4, a consistent comparison with our proposed adaption of IPCA for ARDI, as all three strategies, or combinations of strategies, make use of an IPCA-based algorithm. Alongside the presentation of these algorithms, we propose an explanation of the concept of a fixed point, which is fundamental to IPCA algorithms. Understanding the concept of a fixed point is not only essential for these algorithms, but also crucial for appreciating parameter optimization.
@@ 2.1 ADI with IPCA
Figure 1: Best disk image, according to the SSIM metric, obtained with the IPCA algorithm leveraging ADI only while processing a synthetic dataset (disk #B, cube #3 at a contrast of $10^{-3}$; see Sect. 4). The image on the left displays different apertures placed on (1) the brightest part of the disk, (2) a region of the disk that is rotation invariant, (3) a noise region, and (4) the region near the coronagraphic region where the variation of the speckle field is higher. On the right-hand side, we plot the evolution of the integrated flux computed in these different apertures through the IPCA iterations. The red vertical line indicates the position of the best image among all iterations according to SSIM. Each curve in the figure is normalized by its mean. The x-axis represents each iteration of the IPCA process, starting at rank 1, up to rank 10, with 10 iterations per rank. The gray horizontal line, placed at $y=1$, indicates the position where the flux is equal to its mean.
@@
Figure 2: Same as Fig. 1, but for RDI-only IPCA. The set of reference frames used here is presented in Sect. 4 and is referred to as “optimal.” The x-axis represents each iteration of the IPCA process, for ten iterations using a rank-1 PCA.
In Fig. 1, we show the evolution of the disk estimate through IPCA iterations on an example synthetic dataset. The IPCA is parameterized such that it performs ten iterations per rank, while starting at rank $q=1$ and increasing the rank up to $q=10$. The image on the left provides the best estimation of the injected disk according to the structural similarity index measure (SSIM, Wang et al. 2004). This metric, which we also use in Sect. 4, is particularly adapted for assessing image similarity. It evaluates not only the pixel-wise differences between the images but also considers structural information such as contrast, luminance, and structure on various windows of the two images. Indeed, ranking disk image estimations is a multifactorial problem that often necessitates finding a satisfactory compromise between fewer geometrical biases and fewer noise residuals or errors in the overall restored flux.
@@ 2.2 RDI with IPCA
The implementations of the RDI algorithm found in the scientific literature (Schneider et al. 2009; Soummer et al. 2012; Ren et al. 2018) involves the subtraction of an estimated speckle pattern based on a scaled version of the reference images. We propose an adapted version of the IPCA algorithm that employs an iterative process to optimize the projection of the estimated reference. For a given set of references $L_{r}\in\mathbb{R}^{r,m\times m}$ of $r$ reference images, we subsequently refer to ${V}^{\rm ref}_{q}\in\mathbb{R}^{q,m\times m}$ as the collection of $q$ PCs estimated by PCA on the reference library. We define the operator $\mathcal{H}_{q}^{L_{r}}(Y)$, which, similarly to $\mathcal{H}_{q}(Y)$, returns $\bar{S}_{i+1}$ by projecting the science cube into the principal components, which are now computed on the reference library. We note that $\mathcal{H}_{q}(Y)=\mathcal{H}_{q}^{Y}(Y)$. The subsequent iterative process can then be expressed as:
@@
We display in Fig. 2 the convergence process to find the optimal projection while using IPCA-RDI of rank $q=1$. This example was produced using the same test dataset as in Fig. 1. In this example, we can observe that the flux exhibits a very clear convergence, which is achieved in only a few iterations. This iterative process enables the over-subtraction effect to be mitigated, and significantly improves the results compared to classical PCA. RDI however remains intrinsically sensitive to the stability of the speckle pattern over time, and necessitates the availability of sufficiently correlated reference stars.
@@ 3 Combining angular and reference-star differential imaging (ARDI)
Simultaneously employing both RDI and ADI strategies is expected to help mitigate the individual limitations of each strategy. The combination of RDI with ADI holds the potential of offering several advantages in the field of high-contrast imaging for exoplanet and disk detection around stars. Here, our primary focus is to adapt the IPCA approach, initially designed for ADI, to leverage ARDI. Other advanced algorithms using both ADI and RDI, including IP-based approaches, could also be considered, but are beyond the scope of this paper. In this section we also explore the problem of parameter optimization for IPCA.
@@
where the operation [_;_]__[\_\;;\;\_][ _ ; _ ] represents the concatenation of two collections of images, creating a set of images ($L_{n+r,i}$) with a size of $\mathbb{R}^{n+r,m\times m}$ that contains both the data and reference images. The ratio between the size $r$ of the reference library and the number $n$ of frames in the ADI cube must be chosen carefully for optimal results. We refer to $r/n$ as the reference frame ratio. The proposed algorithm can be classified as semi-supervised, as it simultaneously leverages reference frames for the speckle field (which can be regarded as labeled data), and the actual (unlabeled) data to provide an optimal estimation.
Figure 4: Evolution of the IPCA with ARDI estimate through the iterations, using the same dataset as in Fig. 3. The blue curve relates to the left y-axis and displays the value of the SSIM for each estimate. The orange curve relates to the right y-axis and represents the gradient computed over the full frame, as described in Eq. 4. The $g_{\rm tol}$ has been set to $10^{-2}$ and is represented on the figure by the two gray lines. The black dot indicates the iteration where the gradient curve reaches the $g_{\rm tol}$ and does not exceed it in further steps. The x-axis represents each iteration, where the minor ticks represent a single iteration, and the major ticks represent a rank update.
In Fig. 3, we monitor the evolution of the disk estimate through the IPCA iterations while using the ARDI strategy, similarly to the test presented in Fig. 1 with ADI only. We used identical parametrization (ten ranks, ten iterations per rank), and a reference frame ratio ($r/n$) of one. Conversely to what we observed with ADI only, three out of the four apertures seem to converge when using reference images. Results obtained with ARDI do not exhibit the deformation due to rotation invariant flux (located in aperture #2). However, the background noise in the final image increases throughout the iterations. Consequently, determining where to stop the iterative process involves choosing a good compromise between noise subtraction (non-propagation), filtering, and recovery of extended signals (convergence). A common criterion to assess convergence is setting a stopping criterion based on the gradient of the iterate. In our case, the criterion, based on the disk estimate, reads as follows:
@@
Figure 5: Pearson correlation coefficient of the selected reference frames computed individually and averaged over each frame of the raw ADI data before the injection of simulated disks or planets, for three different selections of references: (left) optimal references selected; (middle) a random sample of each cube’s dedicated references, including the optimal references, in an even proportion; and (right) a random sample of references excluding the optimal references. These coefficients were calculated within an annulus with an outer radius of 60 pixels and inner radius of 30 pixels, and are classified from the most correlated to the least correlated. The text box below each figure displays the mean PCC ($\mu$) and the variance ($\sigma$) of the reference sample relative to its associated data cube.